# Asynchronous Federated Reinforcement Learning with Policy Gradient Updates: Algorithm Design and Convergence Analysis

### v1

@@ 4.3. Key Lemmas
(c) Humanoid-v4
Figure 3. Virtualization of the three tasks in the MuJoCo environments.
Proof.
(16) $-J(\theta_{k+1})~{}$ $\leq~{}-J(\theta_{k})-\langle\nabla J(\theta_{k}),\theta_{k+1}-\theta_{k}\rangle+\frac{L_{g}}{2}\|\theta_{k+1}-\theta_{k}\|^{2}$
If $\|e_{k}\|\leq\frac{1}{2}\|\nabla J(\theta_{k})\|$, we have
(17) $-\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}~{}$ $=~{}-\frac{\|\nabla J(\theta_{k})\|^{2}+\langle\nabla J(\theta_{k}),e_{k}\rangle}{\|d_{k-\delta_{k}}\|}$
$\leq~{}\frac{-\|\nabla J(\theta_{k})\|^{2}+\frac{1}{2}\|\nabla J(\theta_{k})\|^{2}}{\|\nabla J(\theta_{k})\|+\|e_{k}\|}\leq~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|.$
If $\|e_{k}\|\geq\frac{1}{2}\|\nabla J(\theta_{k})\|$, we have
(18) $-\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}\leq~{}\|\nabla J(\theta_{k})\|$
$=~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|+\frac{4}{3}\|\nabla J(\theta_{k})\|\leq~{}-\frac{1}{3}\|\nabla J(\theta_{k})\|+\frac{8}{3}\|e_{k}\|.$
Combine these two conditions, and plugging the result in (16), the lemma can be proved as
(19) $-J(\theta_{k+1})~{}$ $\leq~{}-J(\theta_{k})-\eta_{k}\frac{\langle\nabla J(\theta_{k}),d_{k-\delta_{k}}\rangle}{\|d_{k-\delta_{k}}\|}+\frac{L_{g}}{2}\eta_{k}^{2}$
Next, we construct the relationship between the average concurrency $\bar{\omega}$ and the average delay $\bar{\delta}$ in Lemma 4.7. This lemma gives the boundary (ours) of delays as a corollary of the result in (Koloskova et al., 2022).
Lemma 4.7.
The average delay $\bar{\delta}$ depends on the average concurrency $\bar{\omega}$ as
(20) $\bar{\delta}=\frac{K+1}{K-1+|\mathcal{C}_{K}|}\bar{\omega}\leq\bar{\omega}\leq N.$
Proof.
(21) $\sum_{i=0}^{k}\delta_{i}+\sum_{i\in\mathcal{C}_{k+1}\setminus\{j_{k+1}\}}\delta^{i}_{k+1}~{}$ $=~{}\sum_{i=0}^{k-1}\delta_{i}+\sum_{i\in\mathcal{C}_{k}\setminus\{j_{k}\}}\delta^{i}_{k}+\omega_{k+1}.$
At the zeroth step, there is no delay. Therefore, we have $\delta^{i}_{0}=0$ for all agents. Unrolling the above expression, at the $K$-th step, we have
(22) $\sum_{i=0}^{K-1}\delta_{i}+\sum_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}\delta^{i}_{K}~{}$ $=~{}\sum_{i=0}^{K}\omega_{k+1}~{}=~{}(K+1)\bar{\omega}.$
According to (7) and $|\mathcal{C}_{K}|\geq 2$, we achieve
(23) $\bar{\delta}~{}=~{}\frac{K+1}{K-1+|\mathcal{C}_{K}|}\bar{\omega}\leq\bar{\omega}\leq N.$
In practice, we need to use all resources, and thus, $\bar{\omega}=\omega_{\max}=N$. Since the maximum concurrency is equal to the number of agents $N$, we have the boundary of the average delay as $\bar{\delta}\leq\omega_{\max}\leq N$. ∎
4.4. Proof sketch of Theorem 4.3
Here, we briefly give the main idea of our proof of Theorem 4.3. The complete details and the technical lemmas are provided in Appendix A.
First, we prove the ascent lemma 4.6 based on Lemma 4.4 and the updating rule of AFedPG in Algorithm 1 and 2. We then plug it into (14) in Lemma 4.5 in the expectation form, and get the following inequality
(24) $J^{\star}-\mathbb{E}[J(\theta_{k+1})]~{}\leq$ $~{}(1-\frac{\sqrt{2\mu}\eta_{k}}{3})\big{(}J^{\star}-\mathbb{E}[J(\theta_{k})]\big{)}$
$~{}+\frac{\eta_{k}}{3}\epsilon_{g}+\frac{8}{3}\eta_{k}\mathbb{E}[\|e_{k}\|]+\frac{L_{g}}{2}\eta_{k}^{2}.$
Second, we bound the expected error term $\mathbb{E}[\|e_{K}\|]$ in the ascent lemma with delayed updates. Here we utilize the normalization and look-ahead sampling in Algorithm 1 and 2 to bound the error with an average delay term $\bar{\delta}$ from Lemma 4.7. With suitable learning rates as $\alpha_{k}=(\frac{1}{k+1})^{\frac{4}{5}}$ and $\eta_{k}=\eta_{0}\frac{1}{k+1}$, the upper bound of $\mathbb{E}[\|e_{K}\|]$ is given as follows
(25) $\mathbb{E}[\|e_{K}\|]{\leq}$ $~{}\alpha_{K}\sigma_{g}+c_{1}\sqrt{\alpha_{K}}\sigma_{g}+\frac{9}{4}c_{2}L_{h}\frac{\eta_{K}^{2}}{\alpha_{K}^{2}}+2c_{3}L_{g}\frac{\eta_{K}^{2}}{\alpha_{K}}\bar{\delta},$
where $c_{1}$, $c_{2}$ and $c_{3}$ are constants, and their values are given in Appendix A.2.
Third, unroll the recursion of (24), and then plug the upper bound result (25) into the unrolled form. We now have the following inequality
(26) $J^{\star}$ $-\mathbb{E}[J(\theta_{K})]~{}\leq~{}\frac{\eta_{0}}{3}\epsilon_{g}+\frac{J^{\star}-\mathbb{E}[J(\theta_{0})]}{(K+1)^{2}}+c_{3}\frac{16\eta_{0}^{2}}{3}\frac{L_{g}}{(K+1)^{\frac{7}{5}}}\bar{\delta}$
$+\frac{\eta_{0}^{2}}{2}\frac{L_{g}}{K+1}+\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{4}{5}}}+c_{1}\frac{8\eta_{0}}{3}\frac{\sigma_{g}}{(K+1)^{\frac{2}{5}}}+6c_{2}\eta_{0}^{3}\frac{L_{h}}{(K+1)^{\frac{2}{5}}}.$
With an initial learning rate as $\eta_{0}=\frac{3\mu_{F}}{M_{g}}$, we achieve the global convergence rate in Theorem 4.3.
5. Experiments
5.1. Setup
Environment: To validate the effectiveness of our approach via experiments, we consider three popular MuJoCo environments (Swimmer-v4, Hopper-v4, and Humanoid-v4) (Todorov et al., 2012) with the MIT License. Environmental details are described in Table 2 and the tasks are visualized in Figure 3.
Table 3. Hyperparameter and MLP settings.
@@ 5.1. Setup
MLP $64\times 64$ $256\times 256$ $256\times 256\times 256$ $512\times 512\times 512$
@@
Timesteps ($T$) $2048$ $1024$ $512$ $512$
Iterations ($K$) $1\times 10^{3}$ $2\times 10^{3}$ $3\times 10^{3}$ $1.5\times 10^{4}$
@@
Figure 4. Reward performances of AFedPG ($N=2,4,8$) and PG ($N=1$) on various MoJuCo environments, where $N$ is the number of federated agents. The shadowed areas are confidence intervals with $95\%$ confidence level.
Measurement: All convergence performances are measured over $10$ runs with random seeds from 00 to $9$. The solid lines in Figure 4 and 5 are the averaged results, and the shadowed areas are confidence intervals with the confidence level $95\%$. The lines are smoothed for better visualization.
Implementation: Policies are parameterized by fully connected multi-layer perceptions (MLPs) with settings listed in Table 3. We follow the practical settings in stable-baselines3 (Raffin et al., 2021) to update models with generalized advantage estimation (GAE) ($0.95$) (Schulman et al., 2018) in our implementation. We use PyTorch (Paszke et al., 2019) to implement deep neural networks. All tasks are trained on NVIDIA A100 GPUs with $40$ GB of memory.
Performance metrics: We consider performance metrics as follows:
  1. (1)
  2. (2)
  3. (3)
5.2. Results
Sample complexity: First, to verify the improvement of sample complexity in the first row of Table 1, we evaluate the speedup effects of the number of federated agents $N$. In Figure 4, with different numbers of agents, we test the convergence performances of AFedPG ($N=2,4,8$) and the single agent PG ($N=1$). The x-axis is the number of samples collected by each agent on average, and the y-axis is the reward. In all three tasks, AFedPG beats PG: AFedPG converges faster, has lower variances, and achieves higher final rewards when more agents are involved in collecting trajectories and estimating policy gradients.
(a) Swimmer-v4
(b) Hopper-v4
(c) Humanoid-v4
Figure 5. Global time of AFedPG and FedPG with certain numbers of collected samples on various MoJuCo environments, where $N$ is the number of federated agents. The shadowed areas are confidence intervals with $95\%$ confidence level.
Figure 6. The time complexity speedup of AFedPG compared to FedPG. The x-axis is the heterogeneous ratio of computing power measured by $\frac{t_{\max}}{t_{\min}}$.
Global time complexity: Second, to verify the improvement of global time complexity in the second row of Table 1, we compared the time consumption in the asynchronous and synchronous settings. In Figure 5, we set $N=4,8$ and fix the number of samples collected by all agents. Here, $t_{\max}$ is about $4$ times more than $t_{\min}$. The numbers of total trajectories are $8\times 10^{3}$, $16\times 10^{3}$, and $8\times 10^{4}$ for the Swimmer task, Hopper task, and Humanoid task individually when $N=8$. When $N=4$, the numbers are halved. In all three environments, AFedPG has much lower time consumption compared to the synchronous FedPG.
Effect of computation heterogeneity: At last, we study the effect of the computation heterogeneity among federated agents. The heterogeneity of computing powers is measured by the ratio $\frac{t_{\max}}{t_{\min}}$, and the effect is measured by the speedup, which is the global time of FedPG divided by that of AFedPG. Without loss of generality, we test the performances with one dragger here, which is one agent has $t_{\max}$ time consumption, and all the other agents have $t_{\min}$ time consumption. In Figure 6, the results show that the speedup increases in a nearly linear behavior as the heterogeneity ratio $\frac{t_{\max}}{t_{\min}}$ increases. On the other hand, the more agents that participate in the training process, the higher speedup they achieve. Therefore, AFedPG has the potential to scale up into a very large mobile system, especially when there are some extreme stragglers.
6. Conclusion
In this paper, we proposed AFedPG, a novel asynchronous federated reinforcement learning framework, which updates the global model using policy gradients from $N$ agents. To handle the challenge of lagged policies in the asynchronous setting, we designed a delay-adaptive lookahead technique and used normalized updates to integrate policy gradients. With the above methods, AFedPG achieves speedup for both sample complexity and time complexity. First, our AFedPG method achieves $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ sample complexity at each agent. Compared to the single agent setting (PG) with $\mathcal{O}(\epsilon^{-2.5})$ sample complexity, it enjoys a linear speedup with respect to the number of agents $N$. Second, compared to synchronous FedPG, AFedPG improves the time consumption from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$, where $t_{i}$ denotes the time consumption in each iteration at the agent $i$, and $t_{\max}$ is the largest one. The latter complexity $\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}})$ is always smaller than the former one, and this improvement can be significant in a large-scale federated setting with heterogeneous computing powers ($t_{\max}\gg t_{\min}$). Finally, we empirically verified the performances of AFedPG compared to PG and FedPG in three MuJoCo environments with different $N$. We also demonstrate the improvements with different computing heterogeneity. In summary, AFedPG achieves speedup in terms of both sample complexity and time complexity when it is scaled up to a large mobile system, especially with computing power heterogeneity.
Appendix A Proofs
A.1. Technical Lemmas
Lemma A.1.

### v5

@@ 6.1 Setup
(c) Walker2D-v4
(d) Humanoid-v4
Figure 2: Visualization of the four MuJoCo tasks considered in this paper for experiments.
6.2 Results
Sample complexity improvement: First, to verify the improvement of sample complexity in the first row of Table 1, we evaluate the speedup effects of the number of federated agents $N$. In Figure 3, with different numbers of agents, we test the convergence performances of AFedPG ($N=2,4,8$) and the single agent PG ($N=1$). The x-axis is the number of samples collected by each agent on average, and the y-axis is the reward. In all four MuJoCo tasks, AFedPG beats the single-agent PG: AFedPG converges faster, has lower variances, and achieves higher final rewards when more agents are involved in collecting trajectories and estimating policy gradients. These results confirm the advantage of AFedPG in terms of sample complexity.
(a) Swimmer-v4
(b) Hopper-v4
(c) Walker2D-v4
(d) Humanoid-v4
Figure 3: Reward performances of AFedPG ($N=2,4,8$) and PG ($N=1$) on various MuJoCo environments, where $N$ is the number of federated agents. The solid lines are averaged results over $10$ runs with random seeds from 00 to $9$. The shadowed areas are confidence intervals with $95\%$ confidence level.
(a) Swimmer-v4
(b) Hopper-v4
(c) Walker2D-v4
(d) Humanoid-v4
Figure 4: Global time of AFedPG and FedPG with certain numbers of collected samples on various MuJoCo environments, where $N$ is the number of federated agents. The solid lines are averaged results over $10$ runs. The shadowed areas are confidence intervals with $95\%$ confidence level.
Speedup in global time complexity: Second, to verify the improvement of global time complexity in the second row of Table 1, we compared the time consumption in the asynchronous and synchronous settings. In Figure 4, we set $N=4,~{}8$ and fix the number of samples collected by all agents. Here, $t_{\max}$ is about $4$ times more than $t_{\min}$. The numbers of total samples (trajectories) are $8\times 10^{3}$, $1.6\times 10^{4}$, $8\times 10^{4}$, and $1.2\times 10^{5}$ for the Swimmer-v4 task, the Hopper-v4 task, the Walker2D-v4 task, and the Humanoid-v4 task individually when $N=8$. When $N=4$, the numbers are halved. In all four environments, AFedPG has much lower time consumption compared to the synchronous FedPG, confirming the enhancement in terms of time complexity. Compared to the A3C baseline, AFedPG achieves much higher rewards with less variance.
Ablation study: In Figure 4, we also observe the impact of our delay-adaptive lookahead approach. It is seen that the vanilla scheme without the delay-adaptive lookahead technique does not provide a satisfactory performance, confirming the importance of the proposed approach. We also analyze the effect of computation heterogeneity in Appendix A.3.
7 Discussions
Conclusions: We proposed AFedPG, a novel asynchronous FedRL framework that updates the global model using PGs from multiple agents. To handle the challenge of lagged (heterogeneous) policies in the asynchronous setting, we designed a delay-adaptive lookahead technique and used normalized updates to integrate PGs. We then analytically characterized the convergence bound of AFedPG and showed both global and first-order stationary point convergence guarantees. We also showed that AFedPG achieves a speedup for both sample complexity and time complexity. First, our AFedPG method achieves $\mathcal{O}(\frac{{\epsilon}^{-2.5}}{N})$ sample complexity at each agent for global convergence. Compared to the SOTA result in the single agent setting, i.e. PG, with $\mathcal{O}(\epsilon^{-2.5})$ sample complexity, it enjoys a linear speedup with respect to the number of agents $N$. Second, compared to synchronous FedPG, AFedPG improves the time complexity from $\mathcal{O}(\frac{t_{\max}}{N})$ to $\mathcal{O}({\sum_{i=1}^{N}\frac{1}{t_{i}}})^{-1}$, where $t_{i}$ denotes the time complexity in each iteration at the agent $i$, and $t_{\max}$ is the largest one. The latter complexity $\mathcal{O}({\sum_{i=1}^{N}\frac{1}{t_{i}}})^{-1}$ is always smaller than the former one, and this improvement is significant in a large-scale federated setting with heterogeneous computing powers ($t_{\max}\gg t_{\min}$). Finally, we empirically verified the performances of AFedPG compared to various baselines in four MuJoCo environments with different $N$. We also demonstrated improvements with different computing heterogeneity. It is shown that AFedPG achieves speedup in terms of both sample complexity and time complexity, especially in scenarios with high computing power heterogeneity.
Limitations and future work directions: (i) Although the system is resilient to stragglers, addressing threats posed by adversarial or malicious workers in this setup remains an open problem. While robustness to such attacks has been explored in the synchronous setting (Ganesh et al., 2024a), extending these methods to an asynchronous environment is both a challenging and promising avenue for future research. (ii) It is worth studying whether the asynchronous method is compatible with other methods, e.g., local update, quantization, and low-rank decomposition, to improve communication efficiency in the deep RL setting. (iii) Extending second-order policy optimization methods, such as natural policy gradient methods, which achieve optimal sample complexity in centralized settings (Mondal & Aggarwal, 2024), to an asynchronous federated setup is an important direction for future work. (iv) While this paper focuses on discounted rewards, exploring such results for average reward setup (Bai et al., 2024; Ganesh et al., 2025; 2024b) is an important open direction.
Reproducibility Statement
In this statement, we discuss the efforts that have been made to ensure reproducibility.
For theoretical results, we clearly explain all assumptions in Appendix B.1, and a complete proof of the lemmas and theorems in Appendix B.
For algorithms, we provide the pseudocode in Algorithm 1 and Algorithm 2.
For datasets used in the experiments, we use open source datasets, and describe them in Section 6.1 and Appendix A.2.
Acknowledgments
This work was supported in part by the National Science Foundation (NSF) under grants CPS-2313109 and ITE-2326898, and by the Office of Naval Research (ONR) under grants N00014-22-1-2305 and N00014-23-1-2532.
Appendix A Supplementary Results
In this section, we first compare the time complexity between the synchronous and the asynchronous settings. We then list the experimental settings in Appendix A.2. At last, we have supplementary experiments in Appendix A.3.
A.1 Comparison to the Synchronous Setting
Let $T$ be the computation time during the entire training process to achieve a given number $K$ of cumulative communication rounds of all agents, where $K=\mathcal{O}({\epsilon}^{-2.5})$ for a global convergence according to Theorem 5.1.
In AFedPG, for agent $i$, the number of communication rounds is $\frac{T}{t_{i}}$. For all agents, the total number is $\sum_{i=1}^{N}\frac{T}{t_{i}}$. As $\sum_{i=1}^{N}\frac{T}{t_{i}}=K$, we have $T=\frac{K}{\sum_{i=1}^{N}\frac{1}{t_{i}}}=\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}}{\epsilon}^{-2.5})$ as a harmonic average of all agents.
In FedPG, the number of global communication rounds on the server is $\frac{T}{t_{\max}}$. As $\frac{T}{t_{\max}}=\frac{K}{N}$, we have $T=\frac{Kt_{\max}}{N}=\mathcal{O}(\frac{t_{\max}}{N}{\epsilon}^{-2.5})\geq\mathcal{O}(\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}}{\epsilon}^{-2.5})$ as the harmonic mean is always smaller or equal to the maximum one.
The asynchronous FedPG achieves better time complexity than the synchronous approach regardless of the delay pattern. The advantage is significant when $t_{\max}\gg t_{min}$, which occurs in many practical settings with heterogeneous computation powers across different agents. We illustrate the advantage of AFedPG over synchronous FedPG in Figure 5. As the server only operates one simple summation, without loss of generality, the time consumption at the server-side is negligible.
It is noticeable that we do not make any assumptions or requirements on $t_{\max}$ in the analysis of AFedPG. In the extreme case, the slowest agent does not communicate with the server, and thus, $t_{\max}$ is infinite. In this scenario, the time consumption of AFedPG does not hurt a lot, while the time consumption of FedPG becomes infinite.
Figure 5: Comparison of time consumptions between synchronous and asynchronous approaches. The circled numbers denote the indices of global steps.
A.2 Supplementary Experimental Settings
In Section 6.1, we list the key experimental settings. We list the rest with details in this subsection. The environmental details (the four MuJoCo tasks) are described in Table 2. The policies $\pi_{\theta}$ are parameterized by fully connected multi-layer perceptions (MLPs) with settings listed in Table 3.
Table 2: Detailed descriptions of the four tasks in the MuJoCo environment.
Action Space State Space
MuJoCo Tasks Agent Types Dimension Dimension
Three-link
Swimmer-v4 swimming robot $2$ $8$
Two-dimensional
Hopper-v4 one-legged robot $3$ $11$
Two-dimensional
Walker2D-v4 bipedal robot $6$ $17$
Three-dimensional
Humanoid-v4 bipedal robot $17$ $376$
Table 3: Hyperparameters of AFedPG and the MLP policy parameterization settings.
@@ A.2 Supplementary Experimental Settings
MLP $64\times 64$ $256\times 256$ $512\times 512$ $512\times 512\times 512$
@@
Timesteps ($T$) $2048$ $1024$ $1024$ $512$
Iterations ($K$) $1\times 10^{3}$ $2\times 10^{3}$ $5\times 10^{3}$ $1.5\times 10^{4}$
@@
A.3 Supplementary Experiments
In this subsection, we show three experimental results to study the effect of computation heterogeneity, communication overhead, and reward performances with long runs.
Effect of computation heterogeneity. We study the effect of the computation heterogeneity among federated agents. The heterogeneity of computing powers is measured by the ratio $\frac{t_{\max}}{t_{\min}}$, and the effect is measured by the speedup, which is the global time of FedPG divided by that of AFedPG. Without loss of generality, we test the performances with (1) One straggler, which is one agent has $t_{\min}$ time consumption, and all the other agents have $t_{\max}$ time consumption. This is the scenario with the largest speedup. (2) One leader, which is one agent has $t_{\max}$ and the others have $t_{\min}$. This is the scenario with the smallest speedup.
In Figure 6, the results show that the speedup increases as the heterogeneity ratio $\frac{t_{\max}}{t_{\min}}$ increases. With one leader, as more agents participate, the speedup approximately decreases to a quadratic function w.r.t. the time heterogeneity ratio. With one straggler, the more agents that participate in the training process, the higher speedup they achieve. The overall results indicate that AFedPG has the potential to scale up to a very large RL system, particularly in scenarios with extreme stragglers (The ratio is large: $t_{\max}\gg t_{\min}$).
(a) With one straggler
(b) With one leader
Figure 6: The time complexity speedup of AFedPG compared to FedPG. $N$ is the number of agents. The x-axis is the heterogeneity ratio of computing power measured by $\frac{t_{\max}}{t_{\min}}$. The y-axis is the global time of FedPG divided by that of AFedPG.
Communication overhead analysis. We compare the communication overhead of FedPG and AFedPG in Figure 7. For neural network parameters, we use the standard format float32. The communication overhead is measured by the number of transmitted bytes. The number of federated agents $N$ is set to $8$. The MuJoCo task is Swimmer-v4. Overall, the commutative communication overhead is similar. However, when we dive deep into the agent side and the server side separately in fine-grained time, AFedPG shows advantages on both sides.
On the agent side, Figure 7 (a) shows the cumulative communication bytes during the training process. The cumulative communication overhead is similar in FedPG and AFedPG. In AFedPG, the faster ones, e.g., Agent 8, communicate more, and the slower ones, e.g., Agent 1, communicate less, which is more reasonable than the equal allocation in the synchronous setting. The Agent 1, in fact, commutes with the server during the training process, but it is insignificant in the plot with a log scale.
The agents have heterogeneous resources, but equal allocation could bring too much burden for the slower ones. In AFedPG, it naturally shifts these burdens to the faster ones considering the local resources.
On the server side, we show the downlink communication overhead in a time window in Figure 7 (b). The time window is set as “one global round in the synchronous setting”. At time step 5, it is higher because two agents communicate with the server in a short time period. The total amounts of AFedPG and FedPG are similar. In AFedPG, it is almost evenly distributed during the time span, while in FedPG, the server has a huge burden with a peak. This makes the server in FedPG require huge resources.
(a) On the agent side
(b) Downlink on the server side
Figure 7: The communication overhead of AFedPG compared to FedPG. The number of agents is $N=8$. (a) The cumulative communication overhead on the agent side. (b) The downlink communication overhead in a time window on the server side.
Reward performances with long runs. To further enhance the results in Figure 3, we extend the running time with more samples. Compared to the results in Figure 3, we increase the number of samples by $5$ times in each Mujoco task in Figure 8. The solid lines are averaged results over $5$ runs with random seeds from 00 to $4$. The shadowed areas are confidence intervals with $95\%$ confidence level.
With enough samples, it basically achieves a similar reward performance for different numbers of agents $N$ in AFedPG. However, the more agents engage, the faster it achieves. Notably, the solid line is the average result with $5$ independent runs. With different numbers of agents, the shadowed area has a large overlap. The more overlaps they have, the more runs that have similar reward performances because of the inherent randomness (with different random seeds) in the deep reinforcement learning tasks.
For the Humanoid-v4 task, though in some runs (shadowed area), PG achieves the optimal reward performance, the average (solid line) performance of PG is relatively lower than the others in AFedPG. The reason is that the Humanoid-v4 task has the largest state and action space, which makes the hyperparameter tuning difficult, e.g., learning rates, and brings huge GPU hours. The hyperparameter setting of PG is suboptimal here, as it has no contribution to our main claim. Recall that we aim to use these experiments to verify the speedup effect in AFedPG in Table 1. The suboptimal hyperparameters of PG in the Humanoid-v4 task do not influence the conclusion: The more agents in AFedPG, the faster the optimal reward will be achieved.
@@ A.3 Supplementary Experiments
Figure 8: Reward performances of AFedPG ($N=2,4,8$) and PG ($N=1$) on various MuJoCo environments, where $N$ is the number of federated agents. The solid lines are averaged results over $5$ runs with random seeds from 00 to $4$. The shadowed areas are confidence intervals with $95\%$ confidence level.
Appendix B Theoretical Proofs
In this section, we give the assumptions in Appendix B.1, the technical lemmas in Appendix B.2, our key lemmas in Appendix B.3, the proof of Theorem 5.1 (the global convergence) in Appendix B.4, and the proof of Theorem 5.2 (the FOSP convergence) in Appendix B.5.
B.1 Assumptions
In order to derive the global convergence rates, we make the following standard assumptions (Agarwal et al., 2021a; Ding et al., 2020; Liu et al., 2020; Papini et al., 2018; Xu et al., 2020) on policy gradients, and rewards.
Assumption B.1.
  1. 1.
[…]
  2. 2.
    The score function is $M_{h}$-Lipschitz continuous. In other words, for all $\theta_{i},~{}\theta_{j}\in\mathbb{R}^{d},~{}s\in\mathcal{S},~{}\text{and}~{}a\in\mathcal{A}$, we have
    $\left\|\nabla\log\pi_{\theta_{i}}(a\mid s)-\nabla\log\pi_{\theta_{j}}(a\mid s)\right\|$ $\leq M_{h}\left\|\theta_{i}-\theta_{j}\right\|.$ (12)
  3. 3.
These standard assumptions naturally state that the reward, the first-order, and the second-order derivatives of the score function are not infinite, and they hold with common practical parametrization methods, e.g., softmax policies. With function approximation $\pi_{\theta}$, the approximation error may not be 00 in practice.
Thus, we follow previous works Liu et al. (2020); Agarwal et al. (2021a); Ding et al. (2022); Huang et al. (2020) with an assumption on the expressivity of the policy parameterization class.
Assumption B.2.
(Function approximation) $\exists~{}\epsilon_{{\rm bias}}\geq 0$ s.t. for all $\theta\in\mathbb{R}^{d}$, the transfer error satisfies
$\mathbb{E}[\big{(}A_{\pi_{\theta}}(s,a)-(1-\gamma)u^{\star}(\theta)^{\top}\nabla\log\pi_{\theta}(a\mid s)\big{)}^{2}]\leq\epsilon_{{\rm bias}},$ (13)
It means that the parameterized policy $\pi_{\theta}$ makes the advantage function $A_{\pi_{\theta}}(s,a)$ approximated by the score function $\nabla\log\pi_{\theta}(a\mid s)$ as the features. This assumption is widely used with Fisher-non-degenerate parameterization. $\epsilon_{{\rm bias}}$ can be very small with rich neural network parameterization, and 00 with a soft-max parameterization (Wang et al., 2020).
To achieve the global convergence, we make a standard assumption on the Fisher information matrix (Liu et al., 2020; Ding et al., 2022; Lan et al., 2023b).
Assumption B.3.
(Positive definite) For all $\theta\in\mathbb{R}^{d}$, there exists a constant $\mu_{F}>0$ s.t. the Fisher information matrix $F_{\rho}(\theta)$ induced by the policy $\pi_{\theta}$ and the initial state distribution $\rho$ satisfies
$F_{\rho}(\theta)\succcurlyeq\mu_{F}\cdot I,$ (14)
where $I\in\mathbb{R}^{d\times d}$ is an identity matrix.
For any two symmetric matrices $A$ and $B$ with the same dimension, $A\succcurlyeq B$ denotes that the eigenvalues of $A-B$ are greater or equal to zero.
B.2 Technical Lemmas
Lemma B.4.